OpenStudy (anonymous):

Use the power rule to differentiate the following: g(x) = 1/(x^1/4)

OpenStudy (anonymous):

rewrite equation first. g (x)= x^(-1/4) g' (x)= -1/4 x ^ (-1/4-1) = -1/4 x ^ (-5/4)